OpenStudy (anonymous):

A table is built of two legs, A and B, and a rod of mass-10 kg and length 5m. A small block of 2 kg is placed on the bridge 0,75 m from leg B. Both the rod and the blocks have constant densities(are uniform). You may assume that the legs are very narrow compared to the length of the rod. Find the force with which each leg supports the rod and the block

OpenStudy (anonymous):

The rod is equally supported by each leg so 10/2 = 5kg for each leg but then you have to add the additional weight of the small block. The proportion of 2 kg that each leg holds is inversely proportional to the distance away from the block. So leg B is holding 85% of the block (.85* 2 = 1.7kg) and leg A is holding 15% of the block ( .15 * 2 = .3kg). Now add 5 kg to each for leg B= 6.7kg leg A = 5.3kg